# SpikeCLIP: A Contrastive Language-Image Pretrained Spiking Neural Network

### v3

@@ 3.1 Dual-Stream Architecture
Given the remarkable success of the Transformer architecture (Vaswani et al., 2017), Zhou et al. (2023a) introduced a spiking variant, called Spikingformer, which achieved cutting-edge accuracy across multiple image classification datasets using event-driven spiking computations. Therefore, we chose to employ Spikingformer in building the processing stream for image data. To enable the conversion of spiking outputs into feature vector representations, a linear readout layer was added on the top of the Spikingformer to build a full-fledged image encoder. This layer also uses a set of learnable weights to integrate the spiking signal generated at different time steps (i.e. Time-Dependent Weight). This approach is beyond the rate code solution, emphasizing the significance of the timing of emitted spikes. Lv et al. (2023) have shown that a relatively simple spiking convolutional neural network, bearing a similar architecture to TextCNN (Kim, 2014), can deliver satisfactory accuracy across a range of datasets in both English and Chinese. Considering the shorter text lengths (up to $20$ tokens) in the datasets used in this study compared to theirs, we opted for a less complex architecture, which incorporates Multilayer Perceptron (MLP) to construct a text encoder (the ablation study for other architectural choices for text data processing can be found in Appendix D). Like the image encoder, a similar readout layer is also applied to transform spiking outputs into feature representations.
3.2 Building Block: Integrate-and-Fire Neuron
Many spiking neuron models could be used to build SNNs (Izhikevich, 2004), and we selected the widely-used first-order leaky integrate-and-fire (LIF) neuron as the building block. Analogous to traditional artificial neuron models, LIF neurons compute a weighted sum of inputs that contributes to the membrane potential $U_{t}$ of the neuron at time step $t$. If this sum sufficiently causes the membrane potential to reach a predefined threshold $U_{\rm thr}$ and excites the neuron, the neuron emits a spike $S_{t}$:
@@ 3.2 Building Block: Integrate-and-Fire Neuron
3.3 Conversion of Inputs into Spike Trains
Spiking neural networks only accept spike trains as inputs, and thus data from any modalities must be converted into spike trains for subsequent processing within SNNs. For image data, we first resized all images to a resolution of $224\times 224$ pixels. Each pixel in each channel can take a value within the range of 00 to $255$. These values were then normalized by subtracting the mean from each pixel and subsequently dividing by the standard deviation. To transform the normalized values into spike trains, these values are replicated over $T$ time steps, and Equation (2) is applied to generate spikes. During the application of this Equation for the creation of spike trains, $W$ is consistently set to $1$, $X$ are normalized values, and $T$ is the number of time steps used for training SNNs. Following (Lv et al., 2023), we want to leverage pre-trained word embeddings to enhance the performance of SNNs. In their approach, a Poisson spike train will be generated for each component of a word embedding with a firing rate proportional to its scale. This requires a substantial number of time steps for precise encoding of word embeddings, leading to a significant increase in energy consumption. To reap the benefits of pre-trained word embeddings while simultaneously reducing energy demands, we also employed Equation (2) to transform word embeddings into their corresponding spike trains, with $X$ taking the values of the pre-trained word embeddings. Our empirical results demonstrated that such a conversion works quite well for multimodal modeling with SNNs.
@@ 3.4 Pre-training SpikeCLIP via Knowledge Distillation
Leveraging its dual-stream architecture, SpikeCLIP is capable of generating feature representations for any given images (denoted as $x_{\text{img}}$) or texts (represented as $x_{\text{txt}}$) with the help of the readout layers. Given a sufficient number of image-text pairs, cross-modal feature alignment can be achieved through contrastive learning. However, as previously noted, there is an insufficiency of manually-annotated image-text pairs available for the joint training of both the image and text encoders. To navigate this hurdle, we chose to use the knowledge distillation technique to pre-train our SpikeCLIP. This approach not only enables the feasibility of pre-training for SNNs, but also substantially alleviates the difficulty in training these networks.
For pre-training the image encoder of SpikeCLIP, we used the ImageNet-1k dataset (Russakovsky et al., 2015), which comprises approximately $1.28$ million images and is denoted as $\mathcal{D}_{\text{img}}$. Following Radford et al. (2021), we applied a variety of prompt templates, such as “A photo of a {label}”, over $1,000$ textual labels to generate nearly $116$ thousand sentences (denoted as $\mathcal{D}_{\text{txt}}$). During the pre-training stage, we aimed to align the feature representations generated by SpikeCLIP with those produced by CLIP for both images and texts. This was achieved by using the following loss function:
$\small\small\mathcal{L}_{\text{PT}}=\sum_{x_{\text{img}}\in\mathcal{D}_{\text{img}}}\left(1-\frac{E_{c}(x_{\text{img}})\cdot E_{s}(x_{\text{img}})}{\|E_{c}(x_{\text{img}})\|\|E_{s}(x_{\text{img}})\|}\right)+\sum_{x_{\text{txt}}\in\mathcal{D}_{\text{txt}}}\left(1-\frac{E_{c}(x_{\text{txt}})\cdot E_{s}(x_{\text{txt}})}{\|E_{c}(x_{\text{txt}})\|\|E_{s}(x_{\text{txt}})\|}\right)$ (3)
where $E_{c}(\cdot)$ denotes the feature representation produced by CLIP, and $E_{s}(\cdot)$ denotes the feature vector generated by SpikeCLIP. The pre-training objective is to maximize the cosine similarity between these two representations. Although we employed dual-stream architectures, we did not introduce additional notations to distinguish between the image and text encoders, as their input modalities can clearly differentiate them. During the pre-training phase, only the parameters of SpikeCLIP are updated. This is achieved by back-propagating the errors through the layers to the word embeddings and adjusting the weights using surrogate gradients (Zenke and Vogels, 2021).
@@ 3.5 Fine-tuning Through Dual-loss Function
During the fine-tuning phase on a downstream dataset, the pre-trained SpikeCLIP generates feature representations for an input image and a set of possible labels using its image and text encoders. Note that a prompt template is applied to each label to generate a sentence before it is fed into the spiking text encoder. The feature representation of the image is then used to compute the cosine similarity with each of the textual label representations. These cosine similarity scores are subsequently converted into a probability distribution using a softmax operation. Given an input image $x_{\text{img}}$, such a derived probability distribution, denoted as $y_{\text{img}}$, is compared to its true distribution $t_{\text{img}}$ (i.e., the ground truth label), and used to calculate the cross-entropy loss as follows:
@@
where $n$ is the number of training instances in a downstream dataset. The surrogate gradient method is also used to update the parameters of the spiking encoders. Note that only the weights of the image encoder are fine-tuned, while those of the text encoder remain fixed. This strategy stems from the observation that the number of textual labels in a specific downstream dataset is significantly smaller compared to those explored during the pre-training phase. This strategy not only makes the training process more stable but also enhances the generalizability of the trained models to unseen labels.
@@
$\small\small\mathcal{L}_{\text{FT}}=\mathcal{L}_{\text{KL}}+\lambda\mathcal{L}_{\text{CE}}$ (6)
where the hyper-parameter $\lambda$ governs the relative importance of the regularization term compared with the cross-entropy loss.
@@ 4 Experiments
Table 1: Accuracy achieved by spiking models on CIFAR10 and CIFAR100 datasets. SpikeCLIP surpasses all the baselines except Spikingformer. Note that unimodal ANN models often outperform their multimodal counterparts on classification datasets. For instance, the conventional CLIP achieved an accuracy of $98.45\%$ and $89.70\%$ on CIFAR10 and CIFAR100 respectively, yet it performed worse than ViT (one of the leading ANN unimodal models) by $0.68\%$ on CIFAR10 and $4.5\%$ on CIFAR100. In comparison, the performance gap between SpikeCLIP and Spikingformer is less than that between CLIP and ViT with a difference of $0.515\%$ on average across these two datasets. These experimental results demonstrate that the performance discrepancy between multimodal SNNs and their unimodal counterparts can be narrower than that observed among ANNs.
  Model Type Time Step Accuracy
@@
Hybrid Training (Rathi and Roy, 2020) Unimodal $125$ $92.22$ $67.87$
Diet-SNN (Rathi and Roy, 2020) Unimodal $10/5$ $92.54$ $64.07$
STBP (Wu et al., 2018) Unimodal $12$ $89.83$-- -
STBP NeuNorm (Wu et al., 2018) Unimodal $12$ $90.53$-- -
TSSL-BP (Zhang and Li, 2020) Unimodal $5$ $91.41$-- -
STBP-tdBN (Zheng et al., 2021) Unimodal $4$ $92.92$ $70.86$
TET (Deng et al., 2022) Unimodal $4$ $94.44$ $74.47$
Spikingformer (Zhou et al., 2023a) Unimodal $4$ $\bm{95.95}$ $\bm{80.37}$
SpikeCLIP Multimodal $4$ $\bm{94.48}$ $\bm{77.69}$
We conducted four sets of experiments. Experiment 4.1 is to evaluate the performance of SpikeCLIP compared to existing SNN baselines in image classification tasks; Experiment 4.2 is to assess the robustness of SpikeCLIP and its zero-shot learning capabilities across various image classification benchmark datasets; Experiment 4.3 is to investigate how the sizes and distributions of pre-training datasets impact SpikeCLIP’s performance, as well as to understand the contribution of different loss functions used during the fine-tuning stage; Experiment 4.4 is to compare the theoretical energy consumption of SpikeCLIP with that of its ANN counterparts.
@@ 4.1 Image Classification
We evaluated SpikeCLIP on two well-established CIFAR10 and CIFAR100 image classification datasets (Krizhevsky, 2009) against nine different spiking baselines, including Hybrid Training (Rathi et al., 2020), Diet-SNN (Rathi and Roy, 2020), STBP (Wu et al., 2018), STBP NeuNorm (Wu et al., 2019), TSSL-BP (Zhang and Li, 2020), STBP-tdBN (Zheng et al., 2021), TET (Deng et al., 2022) and Spikingformer (Zhou et al., 2023b). CIFAR10 comprises $10$ categories, whereas CIFAR100 includes $100$ classes (refer to Appendix B for details about the datasets). For each dataset, we trained all models on the training set and then evaluated them on the corresponding test set. We strictly adhered to the standard training and test splits as specified for each dataset.
As we can see from Table 1, SpikeCLIP outperforms all the baselines except Spikingformer. Although Spikingformer achieved higher performance on these two datasets, the performance gap between SpikeCLIP and Spikingformer is relatively small with a difference of $1.47\%$ on CIFAR10 and $2.68\%$ on CIFAR100. It is noteworthy that unimodal ANN models often perform better than their multimodal counterparts on specific datasets. This is because multimodal models also possess the ability for zero-shot transfer to other datasets, albeit at the cost of some performance loss on a specific dataset. For example, the conventional CLIP (Dosovitskiy et al., 2010) achieved an accuracy of $98.45\%$ and $89.70\%$ on CIFAR10 and CIFER100 respectively, falling short of the performance of ViT (Dosovitskiy et al., 2010) (one of the top-performing ANN unimodal models) by $0.68\%$ on CIFAR10 and $4.5\%$ on CIFAR100. In comparison, the performance gap between SpikeCLIP and Spikingformer is less than that between CLIP and ViT with a difference of $0.515\%$ on average across these two datasets. Although SNNs still lag behind ANNs in terms of accuracy yet, our experimental results demonstrate that the performance gap between multimodal SNNs and their unimodal counterparts might be narrower than that observed among traditional ANN models.
@@ 4.2 Zero-shot Results
Figure 3: Zero-shot results on $6$ image classification datasets. SpikeCLIP achieved results comparable to ScratchCLIP with a negligible difference of $0.72\%$ on average. This demonstrates that SNNs can potentially yield results that are on par with their ANN counterparts in multimodal zero-shot tasks.
As far as we know, no previous spiking neural network offers zero-shot capabilities for image classification. A direct comparison between SpikeCLIP and the conventional CLIP would be inappropriate, given that the latter was trained on $400$ million text-image pairs that are not yet publicly accessible. Therefore, we constructed an ANN baseline, termed ScratchCLIP, which mirrors the architecture of SpikeCLIP but uses traditional artificial neurons instead of spiking ones. To evaluate zero-shot learning capabilities, both ScratchCLIP and SpikeCLIP were only pretrained by distilling knowledge from the CLIP without any following fine-tuning on downstream datasets. ScratchCLIP was trained using the standard backpropagation algorithm. Figure 3 illustrates that SpikeCLIP delivers results comparable to ScratchCLIP across six different image classification datasets in a zero-shot setting. In addition to CIFAR10 and CIFAR100 datasets, we used four other datasets for zero-shot learning evaluation: Caltech101 (Fei-Fei et al., 2004), OxfordIIITPet (Parkhi et al., 2012), STL10 (Coates et al., 2011), and Flowers102 (Nilsback and Zisserman, 2008). SpikeCLIP’s performance was marginally inferior to ScratchCLIP, with a negligible difference of $0.72\%$ on average. This suggests that SNNs can potentially yield results that are on par with their ANN counterparts in multimodal zero-shot tasks despite the ease of training ANNs due to the absence of non-differentiability issue.
To evaluate the robustness of SpikeCLIP, we designed two additional, more challenging test settings. The first involved the introduction of noisy labels that are semantically different from those in the original dataset’s label set, yet difficult to differentiate from them. The second involved the replacement of certain portions of the textual labels with semantically equivalent expressions. In the first setting, we expanded the label sets by a factor of two (×$\times 2$), five (×$\times 5$) and eight (×$\times 8$). In the second setting, textual labels were randomly replaced at rates of $20\%$, $40\%$, $80\%$, and $100\%$. For the detailed methods used to extend label sets and replace original labels, please refer to Appendix F. The empirical results reported in Table 2 suggest that SpikeCLIP consistently exhibited a commendable performance even under these two more challenging settings on both CIFAR10 and STL10 datasets, demonstrating its robustness in scenarios involving noisy and previously unseen labels.
Table 2: Accuracy yielded by SpikeCLIP on CIFER10 and STL10 datasets under two challenging settings. In the first, the label sets were enlarged by a factor of two (×$\times 2$), five (×$\times 5$), and eight (×$\times 8$) through the introduction of noisy labels. In the second, textual labels were randomly replaced at rates of $20\%$, $40\%$, $80\%$, and $100\%$ with semantically equivalent expressions. The empirical results demonstrate the robustness of SpikeCLIP in situations characterized by noisy and unseen labels.
@@
Table 3: Empirical results from ablation studies. These experiments were conducted by excluding the pre-training phase and removing the KL-divergence term from the loss function.

### v4

@@ 3.2 Dual-Stream Architecture
Given the remarkable success of the Transformer architecture [38], Zhou et al. [14] introduced a spiking variant, called Spikingformer, which achieved cutting-edge accuracy across multiple image classification datasets using event-driven spiking computations. Therefore, we chose to employ Spikingformer in building the processing stream for image data. To enable the conversion of spiking outputs into feature vector representations, a readout (MLP) layer was added on the top of the Spikingformer to build a full-fledged image encoder. This layer also uses a set of learnable weights to integrate the spiking signal generated at different time steps (i.e. Time-Dependent Weight). This approach is beyond the rate code solution, emphasizing the significance of the timing of emitted spikes. Lv et al. [40] have shown that a relatively simple spiking convolutional neural network, bearing a similar architecture to TextCNN [49], can deliver satisfactory accuracy across a range of datasets in both English and Chinese. Given the shorter text lengths (up to 20 tokens) in the datasets used in this study compared to previous work, we chose a simpler architecture that employs a multi-layer perceptron (MLP) to construct the text encoder. An ablation study of alternative architectural choices for text data processing can be found in Section 4.4. Like the image encoder, a similar readout layer is also applied to transform spiking outputs into feature representations.
3.3 Building Block: Leaky Integrate-and-Fire Neuron
Various spiking neuron models can be used to construct SNNs, and we chose the widely-used first-order leaky integrate-and-fire (LIF) neuron [50] as the foundational building block. Analogous to traditional artificial neuron models, LIF neurons compute a weighted sum of inputs that contributes to the membrane potential $U_{t}$ of the neuron at time step $t$. If this sum sufficiently causes the membrane potential to reach a predefined threshold $U_{\rm thr}$ and excites the neuron, the neuron emits a spike $S_{t}$:
@@ 3.3 Building Block: Leaky Integrate-and-Fire Neuron
3.4 Converting Inputs into Spike Trains
Spiking neural networks accept spike trains as inputs, and thus data from any modalities must be converted into spike trains for subsequent processing within SNNs.
For image data, we initially resized all images to a resolution of $224\times 224$ pixels. Each pixel in every channel can take a value within the range of 00 to $255$. These pixel values were then normalized by subtracting the mean of each pixel and dividing by its standard deviation. To convert the image data into spike trains, we employed direct encoding, a type of temporal encoding, by replicating the normalized pixel values for $T$ time steps. In this process, the pixel intensity information is not explicitly transformed into spike timing but is instead maintained as a constant input over time. This sustained input influences the membrane potential dynamics of spiking neurons, which in turn governs the temporal evolution of spike generation. As a result, while direct encoding does not directly map pixel intensity to spike timing, it still falls under the category of temporal encoding since it leverages the temporal dimension to propagate information rather than encoding it in a single discrete step. This procedure transforms the normalized pixel values into spike trains, with the corresponding spikes generated according to Equation (1). During the application of this Equation for the creation of spike trains, $W$ is consistently set to $1$, $X$ are normalized values, and $T$ is the number of time steps used for training SNNs.
For text data, we adopt the approach of [40], utilizing pre-trained word embeddings to enhance SNN performance. In this method, a Poisson spike train is generated for each component of a word embedding, with the firing rate proportional to its value. However, this approach requires a large number of time steps to accurately encode word embeddings, leading to increased energy consumption. Similarly, we use Equation (1) to directly transform the word embeddings after $T$ repetition into corresponding spike trains, where $X$ represents the values of the pre-trained embeddings. Our empirical results show that this conversion is effective for multimodal modeling with SNNs.
@@ 3.5 Pre-training SpikeCLIP via Knowledge Distillation
Leveraging the dual-stream architecture, SpikeCLIP is capable of generating feature representations for any given images (denoted as $x_{\text{img}}$) or texts (represented as $x_{\text{txt}}$) with the help of the readout layers. Given a sufficient number of image-text pairs, cross-modal feature alignment can be achieved through contrastive learning. However, as previously noted, there is an insufficiency of manually annotated image-text pairs available for the joint training of both the image and text encoders. To navigate this hurdle, we chose to use the knowledge distillation technique to pre-train our SpikeCLIP. This approach not only enables the feasibility of pre-training for SNNs, but also substantially alleviates the difficulty in training these networks.
For pre-training the image encoder of SpikeCLIP, we used the ImageNet-1k dataset [51], which comprises approximately $1.28$ million images and is denoted as $\mathcal{D}_{\text{img}}$. Following Radford et al. [19], we applied a variety of prompt templates, such as “A photo of a {label}”, over $1,000$ textual labels to generate nearly $116$ thousand sentences (denoted as $\mathcal{D}_{\text{txt}}$). During the pre-training stage, we aimed to align the feature representations generated by SpikeCLIP with those produced by CLIP for both images and texts. This was achieved by using the following loss function:
$\small\begin{split}\small\mathcal{L}_{\text{Pre-train}}\!=\!&\sum_{x_{\text{img}}\in\mathcal{D}_{\text{img}}}\left(1\!-\!\frac{E_{c}(x_{\text{img}})\cdot E_{s}(x_{\text{img}})}{\|E_{c}(x_{\text{img}})\|\|E_{s}(x_{\text{img}})\|}\right)\!\\ &+\!\sum_{x_{\text{txt}}\in\mathcal{D}_{\text{txt}}}\left(1\!-\!\frac{E_{c}(x_{\text{txt}})\cdot E_{s}(x_{\text{txt}})}{\|E_{c}(x_{\text{txt}})\|\|E_{s}(x_{\text{txt}})\|}\right),\end{split}$ (3)
where $E_{c}(\cdot)$ denotes the feature representation produced by CLIP, and $E_{s}(\cdot)$ denotes the feature vector generated by SpikeCLIP. The pre-training objective aims to maximize the cosine similarity between the two representations. While we used dual-stream architectures, we did not introduce separate notations for the image and text encoders, as the input modalities inherently distinguish them. During pre-training, only the parameters of SpikeCLIP are updated by backpropagating errors through the layers to the word embeddings, with weight adjustments made using surrogate gradients.
@@ 3.6 Fine-tuning Through Dual-loss Function
During the fine-tuning phase on a downstream dataset, the pre-trained SpikeCLIP generates feature representations for an input image and a set of possible labels using its image and text encoders. A prompt template is applied to each label to generate a sentence, which is then input into the spiking text encoder. The image feature representation is subsequently used to compute the cosine similarity with each of the textual label representations. These cosine similarity scores are subsequently converted into a probability distribution using a softmax operation. Given an input image $x_{\text{img}}$, such a derived probability distribution, denoted as $y_{\text{img}}$, is compared to its true distribution $t_{\text{img}}$ (i.e., the ground truth label), and used to calculate the cross-entropy loss as follows:
@@
where $n$ is the number of training instances in a downstream dataset. Notably, only the weights of the image encoder are fine-tuned, while the text encoder weights remain fixed. This strategy stems from the observation that the number of textual labels in a specific downstream dataset is significantly smaller compared to those explored during the pre-training phase. This strategy not only makes the training process more stable but also enhances the generalizability of the trained models to unseen labels.
@@
$\small\small\mathcal{L}_{\text{FT}}=\lambda_{1}\mathcal{L}_{\text{KL}}+\lambda_{2}\mathcal{L}_{\text{CE}}$ (6)
where the hyper-parameter $\lambda_{1}$ and $\lambda_{2}$ govern the relative importance of the regularization term compared with the cross-entropy loss.
@@ 4 Experiments
Table 1: Accuracy achieved on CIFAR10 and CIFAR100 datasets. “Spike” denotes whether the model is a SNN. We report the modality type in the “Type” column. The best and second-best results are highlighted in bold and underlined formats, respectively.
  Model Spike Param. (M) Type Time Step Accuracy (%)
@@
ViT [52] $86.39$ Unimodal $\bm{99.13}$ $\bm{94.20}$
Hybrid Training [53] $9.27$ Unimodal $125$ $92.22$ $67.87$
Diet-SNN [53] $0.27$ Unimodal $10/5$ $92.54$ $64.07$
STBP [54] $17.54$ Unimodal $12$ $89.83$-- -
STBP NeuNorm [54] $17.54$ Unimodal $12$ $90.53$-- -
TSSL-BP [55] $17.54$ Unimodal $5$ $91.41$-- -
STBP-tdBN [56] $12.63$ Unimodal $4$ $92.92$ $70.86$
TET [57] $12.63$ Unimodal $4$ $94.44$ $74.47$
Spikingformer [14] $9.32$ Unimodal $4$ $\underline{95.95}$ $\underline{80.37}$
CLIP [19] $149.60$ Multimodal $\bm{98.45}$ $\bm{89.70}$
SpikeCLIP (Ours) $56.87$ Multimodal $4$ $\underline{94.48}$ $\underline{77.69}$
We conducted four sets of experiments. First, we evaluate the performance of SpikeCLIP against existing ANN and SNN baselines in image classification tasks. Second, we assess the robustness of SpikeCLIP and its zero-shot learning capabilities across various image classification benchmark datasets. Third, we investigate the impact of each component on SpikeCLIP’s performance and examine the choice of model architecture and key hyperparameters. Finally, we compare the theoretical computing energy consumption of SpikeCLIP with that of its ANN counterparts. For detailed implementation, datasets, and hyper-parameters, please refer to B, C, D, and F.
@@ 4.1 Image Classification
We evaluated SpikeCLIP on two well-established CIFAR10 and CIFAR100 image classification datasets [58] against two ANN baselines, i.e., ViT [19] and CLIP [19], and nine different spiking baselines, including Hybrid Training [27], Diet-SNN [53], STBP [54], STBP NeuNorm [59], TSSL-BP [55], STBP-tdBN [56], TET [57], TEBN [60] and Spikingformer [15]. For each dataset, we trained all models on the training set and then evaluated them on the corresponding test set. We strictly adhered to the standard training and test splits as specified for each dataset.
As shown in Table 1, SpikeCLIP outperforms all SNN baselines, except for the unimodal Spikingformer. Although Spikingformer achieved higher performance on these two datasets, the performance gap between SpikeCLIP and Spikingformer is relatively small with a difference of $1.47\%$ on CIFAR10 and $2.68\%$ on CIFAR100. In general, unimodal ANN models tend to perform better than their multimodal counterparts on specific datasets. This is because multimodal models also possess the ability for zero-shot transfer to other datasets, albeit at the cost of some performance loss on a specific dataset. For example, the conventional CLIP [19] achieved an accuracy of $98.45\%$ and $89.70\%$ on CIFAR10 and CIFER100 respectively, falling short of the performance of ViT [52] (one of the top-performing ANN unimodal models) by $0.68\%$ on CIFAR10 and $4.5\%$ on CIFAR100. In comparison, the performance gap between SpikeCLIP and Spikingformer is less than that between CLIP and ViT with a difference of $0.515\%$ on average across these two datasets. These experimental results demonstrate that the performance discrepancy between multimodal SNNs and their unimodal counterparts can be narrower than that observed among ANNs.
@@ 4.2 Zero-shot Results
Figure 5: Zero-shot results on $6$ image classification datasets. SpikeCLIP achieved results comparable to ScratchCLIP, with a negligible average difference of only $0.72\%$.
To the best of our knowledge, no previous spiking neural network has demonstrated zero-shot capabilities for image classification. A direct comparison between SpikeCLIP and the conventional CLIP would be inappropriate, given that the latter was trained on $400$ million text-image pairs that are not yet publicly accessible. Therefore, we constructed an ANN baseline, termed ScratchCLIP, which mirrors the architecture of SpikeCLIP but uses traditional artificial neurons instead of spiking ones.
To evaluate zero-shot learning capabilities, both ScratchCLIP and SpikeCLIP were only pre-trained by distilling knowledge from the CLIP without any following fine-tuning on downstream datasets. ScratchCLIP was trained using the standard backpropagation algorithm. In addition to the CIFAR10 and CIFAR100 datasets, we follow the original CLIP setup and use 22 other datasets for zero-shot learning evaluation, including Caltech101 [61], OxfordIIITPet [62], STL10 [63], Flowers102 [64], among others. Selected results are shown in Figure 5, with detailed outcomes provided in E. Figure 5 illustrates that SpikeCLIP delivers results comparable to ScratchCLIP across $6$ different image classification datasets in a zero-shot setting. SpikeCLIP’s performance was marginally inferior to ScratchCLIP, with a negligible difference of $0.72\%$ on average. This suggests that SNNs can potentially yield results that are on par with their ANN counterparts in multimodal zero-shot tasks despite the ease of training ANNs due to the absence of non-differentiability issues.
Table 2: SpikeCLIP’s accuracy on the CIFER10 and STL10 datasets under two challenging settings: first, label sets were enlarged by factors of $2$, $5$, and $8$ through noisy labels; second, textual labels were randomly replaced at rates of $20\%$, $40\%$, $80\%$, and $100\%$ with semantically equivalent expressions.
@@
To assess the robustness of SpikeCLIP, we designed two additional, more challenging test scenarios. The first involved the introduction of noisy labels that are semantically different from those in the original dataset’s label set, yet difficult to differentiate from them. The second involved the replacement of certain portions of the textual labels with semantically equivalent expressions. In the first setting, we expanded the label sets by a factor of two (×$\times 2$), five (×$\times 5$) and eight (×$\times 8$). In the second setting, textual labels were randomly replaced at rates of $20\%$, $40\%$, $80\%$, and $100\%$. For the detailed methods used to extend label sets and replace original labels, please refer to D. The empirical results presented in Table 2 indicate that SpikeCLIP consistently performed well under both challenging settings on the CIFAR10 and STL10 datasets, highlighting its robustness in handling noisy labels and previously unseen label variations.
4.3 Ablation Study
Firstly, to evaluate the impact of knowledge distillation during pre-training and the introduction of the KL-divergence term during fine-tuning on SpikeCLIP’s performance, we conducted a series of ablation studies across six different datasets. These studies involved eliminating the pre-training stage and removing the KL-divergence term from the loss function.
Table 3: Empirical results from ablation studies. These experiments were conducted by excluding the pre-training phase or fine-tuning phase (data from Figure 5), and removing the KL-divergence term from the loss function.